OpenStudy (anonymous):

Factor completely: 9x2 + 42x + 49 (3x + 7)(3x − 7) (9x − 7)(9x − 7) (3x + 7)(3x + 7) (9x + 7)(9x + 7)

OpenStudy (anonymous):

\[9x^2 + 42x + 49\] so the first term is: \(9x^2 \) Then you know: \(3 \times 3 = 9\) \[x \times x = x^2 \] so: \[(3x) \times (3x) = 9x^2 \]

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

Therefore, choices B and D are right away crossed out as a correct possible answer choice.

OpenStudy (anonymous):

notreally

OpenStudy (mertsj):

Use FOIL on each one of your answers and see which one works.

OpenStudy (anonymous):

yes ^ that also works, it is like working backwards if you do not really understand what i was talking about.